# Can LLMs Interpret and Leverage Structured Linguistic Representations? A Case Study with AMRs

### v1

@@ Abstract
This paper evaluates the ability of Large Language Models (LLMs) to leverage contextual information in the form of structured linguistic representations. Specifically, we examine the impact of encoding both short and long contexts using Abstract Meaning Representation (AMR) structures across a diverse set of language tasks. We perform our analysis using 8-bit quantized and instruction-tuned versions of Llama 3.1 (8B), Phi-3, and Mistral 7B. Our results indicate that, for tasks involving short contexts, augmenting the prompt with the AMR of the original language context often degrades the performance of the underlying LLM. However, for tasks that involve long contexts, such as dialogue summarization in the SAMSum dataset, this enhancement improves LLM performance, for example, by increasing the zero-shot cosine similarity score of Llama 3.1 from 66.2% to 76%. This improvement is more evident in the newer and larger LLMs, but does not extend to the older or smaller ones. In addition, we observe that LLMs can effectively reconstruct the original text from a linearized AMR, achieving a cosine similarity of 81.3% in the best-case scenario.
@@ 1 Introduction
2. We demonstrate that while AMR integration tends to degrade performance in short-context tasks, it can enhance performance in long-context tasks, particularly when using larger and more recently developed LLMs. Additionally, in some downstream tasks, LLMs can work exclusively with linearized AMRs (AMR-only prompting) to achieve reasonable performance with few-shot prompting.
3. We show that LLMs can effectively reconstruct the original context from linearized AMRs, achieving 81.3% cosine similarity in the best-case scenario, highlighting their ability to understand context from AMRs.
@@ 5.1 LDC2020T02
Fig. 9 illustrates that Llama3.1 outperforms Phi3 and Mistral in regenerating the original contexts from the linearized LDC2020T02 AMRs using few-shot prompting. While Phi3 achieves the best zero-shot regeneration cosine similarity score (74.3%), it becomes the worst performer with few-shot prompting. With 5-shot prompting, Llama3.1 attains a cosine similarity score of 81.3%. Using more than 3 in-context learning examples yields diminishing returns, as indicated by the marginal improvement when transitioning from 3-shot to 5-shot prompting. Detailed results for this task, including additional regeneration metrics, are provided in Table 2 in the Appendix.
@@ 5.2 SAMSum
Fig. 10 illustrates how Llama3.1’s zero-shot summarization cosine similarity score improves from 66.2% to 76% with AMR-augmented prompting, compared to context-only prompting. Additionally, AMR-only prompting achieves a higher 3-shot cosine similarity score (78.9%) than context-only prompting (78.1%). This provides strong evidence that AMRs can enhance the ability of LLMs, particularly larger and newer models, to infer information from long contexts, especially long dialogues, and retain key pieces of information. However, it should be noted that increasing the number of few-shot examples beyond 3 yields diminishing returns with both context-only and AMR-augmented prompting, while it deteriorates performance with AMR-only prompting. While Llama3.1’s performance improved with AMRs, Phi3 and Mistral did not exhibit similar improvements (Table 4). Mistral was the best performing model for this task, despite being worse than Llama3.1 in all the other tasks.
Rank-32 LoRA fine-tuning of Llama3.1 achieved a cosine similarity score of 75.1% with a context-only input prompt, which increased to 76% with an AMR-augmented input prompt. This indicates that, although LoRA fine-tuning of Llama3.1 was not any better than few-shot prompting for SAMSum summarization, it benefited from AMR augmentation.
Figure 10: Llama3.1 SAMSum summarization Cosine Similarity.
@@
Fig. 11 illustrates that Llama3.1’s single-hop reasoning F1-score declines with AMR-augmented prompting compared to context-only prompting (e.g., from 59.1% to 52.3% with 3-shot prompting). However, both Llama3.1 (Fig. 11) and Mistral (Table 3) demonstrate a remarkable understanding of AMRs in 3-shot single-hop reasoning, as indicated by how closely their 3-shot AMR-only performance approaches the best possible 3-shot performance for this task. Including over 3 examples in the prompt results in only marginal performance improvements with context-only prompting. However, with AMR-augmented and AMR-only prompting, performance deteriorates as the number of few-shot examples increases to 5. For instance, Llama3.1 achieved a 47.7% F1 score with AMR-only 3-shot prompting, which dropped significantly to 25.6% with AMR-only 5-shot prompting. Detailed results for this task, including results from other models, are provided in Table 3 in the Appendix.
Figure 11: Llama3.1 SQuAD 2.0 QA F1-score.
@@ 5.4 HotpotQA
Figure 12: Llama3.1 2-hop HotpotQA F1-score.
@@ 5.5 SNLI
Figure 13: Phi3 SNLI Macro F1-score.
@@ 5.6 DocNLI
The DocNLI experiment was conducted on approximately 13k examples from the DocNLI test split, which had an 8:5 label imbalance in favor of contradiction. This experiment aimed to further validate observations from prior experiments regarding the utility of AMRs in long-context tasks. Fig. 14 illustrates that AMR-only prompting achieves the highest zero-shot macro F1 score of 20%. However, with 3 few-shot examples, context-only prompting achieves a macro F1 score of 51%, outperforming the other prompting methods. Both AMR-only and context-only prompting result in a performance decline with 5 few-shot examples, while the AMR-augmented prompting macro F1 score increases to 50%. These results are not only inconsistent with the SNLI experiment outcome but also deviate from the trends observed in other long-context tasks.
Figure 14: Llama3.1 DocNLI Macro F1-score.
@@ 6 Conclusion
In this study, we investigated the ability of LLMs to interpret and leverage AMRs. Our findings demonstrate that AMR-augmented prompting, where the AMR of the context is included alongside the original context, significantly improves zero-shot long dialogue summarization and sentence-level natural language inference for Llama3.1, the largest and most recently developed model examined in this work. While AMRs did not enhance performance in other tasks, there was still some evidence that LLMs can extract meaningful information from AMRs. This was particularly evident in some of the AMR-only experiments, where the models achieved reasonable performance using AMRs alone. Additionally, we conducted rank-32 LoRA fine-tuning of Llama3.1 on the SAMSum summarization task, which produced results better than zero-shot but not as effective as few-shot prompting for both context-only and AMR-augmented approaches.
@@ Appendix A Appendix
Llama3.1 0 Context-only 54.8 66.0
3 Context-only 59.1 67.8
5 Context-only 59.4 68.1
0 AMR-augmented 48.8 63.1
3 AMR-augmented 52.3 62.3
5 AMR-augmented 48.9 59.5
0 AMR-only 17.8 38.5
3 AMR-only 47.7 60.3
5 AMR-only 25.6 45.0
Phi3 0 Context-only 45.9 56.3
3 Context-only 47.0 57.5
0 AMR-augmented 38.2 46.2
3 AMR-augmented 40.4 53.2
0 AMR-only 20.1 40.0
3 AMR-only 21.9 42.1
Mistral 0 Context-only 35.4 46.8
3 Context-only 50.8 61.1
0 AMR-augmented 27.4 37.4
3 AMR-augmented 48.7 60.2
0 AMR-only 16.7 35.0
3 AMR-only 40.7 55.4
@@
Llama3.1 0 Context-only 66.2 29 8.2 21.3 2
3 Context-only 78.1 42.4 17 33.6 7.3
5 Context-only 78 42.9 17 34.4 7.4
LoRA Context-only 75.1 41.1 16 32.8 6.6
0 AMR-augmented 76 36.7 12.3 27.9 4.3
3 AMR-augmented 79 41.1 17.2 31.9 8.2
5 AMR-augmented 79.1 43.3 18.2 33.9 9.3
LoRA AMR-augmented 76 42.7 17.1 34.3 6.8
0 AMR-only 60.2 27.8 5.3 19.3 1.1
3 AMR-only 78.9 41.1 17.2 31.9 8.2
5 AMR-only 67.2 31.7 7.8 22.2 2.3
Phi3 0 Context-only 71 27.8 5.8 20.3 0.9
3 Context-only 72.3 30.7 8.2 23.2 1.6
0 AMR-augmented 70.2 26.4 5.3 18.8 0.7
3 AMR-augmented 70.1 25.1 5.6 18.3 0.8
0 AMR-only 58.7 21.9 2.3 13.8 0.2
3 AMR-only 55.1 18.1 1.7 11.8 0.1
Mistral 0 Context-only 76.3 36.1 12 27.3 4.6
3 Context-only 79.7 47.2 22.1 37.7 12.1
0 AMR-augmented 76.6 38 14.1 28.8 5.8
3 AMR-augmented 76.6 44.8 20.4 35.5 11.1
0 AMR-only 66.1 30 6.3 20.7 1.2
3 AMR-only 66 33.7 8 24.6 2.1
@@
Llama3.1 0 Context-only 46.9 8
3 Context-only 55.6 52
0 AMR-augmented 42.2 15
3 AMR-augmented 56.3 54
0 AMR-only 37.3 12
3 AMR-only 43.6 39
Phi3 0 Context-only 77.4 27
3 Context-only 81.6 80
5 Context-only 83.2 82
0 AMR-augmented 66.7 39
3 AMR-augmented 79.2 79
5 AMR-augmented 75.7 76
0 AMR-only 49.5 25
3 AMR-only 52.1 52
5 AMR-only 54.5 55
Mistral 0 Context-only 48.4 50
3 Context-only 52.4 54
0 AMR-augmented 48.5 50
3 AMR-augmented 33.6 20
0 AMR-only 32.7 20
3 AMR-only 32.8 12

### v4

@@ Abstract
This paper evaluates the ability of Large Language Models (LLMs) to leverage contextual information in the form of structured linguistic representations. Specifically, we examine the impact of encoding both short and long contexts using Abstract Meaning Representation (AMR) structures across a diverse set of language tasks. We perform our analysis using 8-bit quantized and instruction-tuned versions of Llama 3.1 (8B), Phi-3, and Mistral 7B. Our results indicate that, for tasks involving short contexts, augmenting the prompt with the AMR of the original language context often degrades the performance of the underlying LLM. However, for tasks that involve long contexts, such as dialogue summarization in the SAMSum dataset, this enhancement improves LLM performance, for example, by increasing the zero-shot cosine similarity score of Llama 3.1 from 66% to 76%. This improvement is more evident in the newer and larger LLMs, but does not extend to the older or smaller ones. In addition, we observe that LLMs can effectively reconstruct the original text from a linearized AMR, achieving a cosine similarity of 81% in the best-case scenario.
@@ 1 Introduction
2. We demonstrate that while AMR integration tends to degrade performance in short-context tasks, it can enhance performance in long-context tasks, an observation that is more profound when using larger and more recently developed LLMs. Additionally, in some downstream tasks, LLMs can work exclusively with linearized AMRs (AMR-only prompting) to achieve reasonable performance with few-shot prompting.
3. We show that LLMs can effectively reconstruct the original context from linearized AMRs, achieving 81% cosine similarity in the best-case scenario, highlighting their ability to understand context from AMRs.
@@ 5 Results
For all the zero-shot experiments except AMR-to-text, 5 different transformer seeds were used to compute 90% confidence intervals. Similar performance was recorded across all seeds for zero-shot experiments, resulting in narrow confidence intervals. For all the few-shot experiments except AMR-to-text, 5 distinct sets of examples were curated from the training set to determine 90% confidence intervals. All the scores presented in the results and tables are averaged across different seeds and sets, if applicable.
@@ 5.1 LDC2020T02
Fig. 9 illustrates that Llama3.1 outperforms Phi3 and Mistral in regenerating the original contexts from the linearized LDC2020T02 AMRs using few-shot prompting. While Phi3 achieves the best zero-shot regeneration cosine similarity score (74%), it becomes the worst performer with few-shot prompting. With 5-shot prompting, Llama3.1 attains a cosine similarity score of 81%. Using more than 3 in-context learning examples yields diminishing returns, as indicated by the marginal improvement when transitioning from 3-shot to 5-shot prompting. Detailed results for this task, including additional regeneration metrics, are provided in Table 2 in the Appendix.
@@ 5.2 SAMSum
Figure 10 illustrates how Llama3.1’s zero-shot summarization cosine similarity score improves from 66% to 76% with AMR-augmented prompting, compared to context-only prompting. In the few-shot setting, although the confidence intervals for context-only and AMR-augmented prompting overlap, both the average and maximum cosine similarity scores are higher for AMR-augmented prompting than for other prompting strategies. This provides strong evidence that AMRs can enhance the ability of LLMs, particularly larger and more recent models, to infer information from long contexts, especially lengthy dialogues, and to retain key pieces of information. However, increasing the number of few-shot examples beyond 3 yields diminishing returns for both context-only and AMR-augmented prompting, and it even degrades performance with AMR-only prompting.
While Llama3.1’s performance improved with AMR augmentation, Phi3 and Mistral did not exhibit similar gains (Table 4). Mistral was the best performing model for this task by a small margin, despite being worse than Llama3.1 in all the other tasks. Rank-32 LoRA fine-tuning of Llama3.1 achieved a cosine similarity score of 75% with a context-only input prompt, which increased to 76% with an AMR-augmented input prompt. This indicates that, although LoRA fine-tuning of Llama3.1 was not any better than few-shot prompting for SAMSum summarization, it benefited marginally from AMR augmentation.
Figure 10: Llama3.1 SAMSum summarization Cosine Similarity with 90% confidence intervals.
@@
Fig. 11 illustrates that Llama3.1’s single-hop reasoning F1-score declines with AMR-augmented prompting compared to context-only prompting (e.g., from 59% to 52% with 3-shot prompting). However, both Llama3.1 (Fig. 11) and Mistral (Table 3) demonstrate a remarkable understanding of AMRs in 3-shot single-hop reasoning, as indicated by how closely their 3-shot AMR-only performance approaches the best possible 3-shot performance for this task. Including over 3 examples in the prompt results in only marginal performance improvements with context-only prompting. However, with AMR-augmented and AMR-only prompting, performance deteriorates as the number of few-shot examples increases to 5. For instance, Llama3.1 achieved a 48% F1 score with AMR-only 3-shot prompting, which dropped significantly to 26% with AMR-only 5-shot prompting. Detailed results for this task, including results from other models, are provided in Table 3 in the Appendix.
Figure 11: Llama3.1 SQuAD 2.0 QA F1-score with 90% confidence intervals.
@@ 5.4 HotpotQA
Figure 12: Llama3.1 2-hop HotpotQA F1-score with 90% confidence intervals.
@@ 5.5 SNLI
Figure 13: Phi3 SNLI Macro F1-score with 90% confidence intervals.
@@ 5.6 DocNLI
The DocNLI experiment was conducted on approximately 13k examples from the DocNLI test split, which had an 8:5 label imbalance in favor of contradiction. This experiment aimed to further validate observations from prior experiments regarding the utility of AMRs in long-context tasks. Fig. 14 illustrates that AMR-only prompting achieves the highest zero-shot macro F1 score of 20%. However, with 3 few-shot examples, context-only prompting achieves a macro F1 score of 51%, outperforming the other prompting methods. Both AMR-only and context-only prompting result in a performance decline with 5 few-shot examples, while the AMR-augmented prompting macro F1 score increases. These results are not only inconsistent with the SNLI experiment outcome but also deviate from the trends observed in other long-context tasks.
Figure 14: Llama3.1 DocNLI Macro F1-score with 90% confidence intervals.
@@ 6 Conclusion
In this study, we investigated the ability of LLMs to interpret and leverage AMRs. Our findings demonstrate that AMR-augmented prompting, where the AMR of the context is included alongside the original context, significantly improves zero-shot long-dialogue summarization for Llama3.1 and noticeably improves sentence-level natural language inference for Phi3. The improvement from AMR-augmented prompting is more significant in SAMSum summarization with Llama3.1, the largest and most recently developed model among those used in this study, compared to other models and tasks. While AMRs did not enhance performance in other tasks, there was still some evidence that LLMs can extract meaningful information from AMRs. This was particularly evident in some of the AMR-only experiments, where the models achieved reasonable performance using AMRs alone. Additionally, we conducted rank-32 LoRA fine-tuning of Llama3.1 on the SAMSum summarization task, which produced results better than zero-shot but not as effective as few-shot prompting for both context-only and AMR-augmented approaches.
@@ Appendix A Appendix
Llama3.1 0 Context-only 55 66
3 Context-only 59 68
5 Context-only 59 68
0 AMR-augmented 49 63
3 AMR-augmented 52 62
5 AMR-augmented 49 60
0 AMR-only 18 38
3 AMR-only 48 60
5 AMR-only 26 45
Phi3 0 Context-only 46 56
3 Context-only 47 58
0 AMR-augmented 38 46
3 AMR-augmented 40 53
0 AMR-only 20 40
3 AMR-only 22 42
Mistral 0 Context-only 35 47
3 Context-only 51 61
0 AMR-augmented 27 37
3 AMR-augmented 49 60
0 AMR-only 17 35
3 AMR-only 41 55
@@
Llama3.1 0 Context-only 66 29 8 21 2
3 Context-only 78 42 17 34 7
5 Context-only 78 43 17 34 7
LoRA Context-only 75 41 16 33 7
0 AMR-augmented 76 37 12 28 4
3 AMR-augmented 79 41 17 32 8
5 AMR-augmented 79 43 18 34 9
LoRA AMR-augmented 76 43 17 34 7
0 AMR-only 60 28 5 19 1
3 AMR-only 79 41 17 32 8
5 AMR-only 67 32 8 22 2
Phi3 0 Context-only 71 28 6 20 1
3 Context-only 72 31 8 23 2
0 AMR-augmented 70 26 5 19 1
3 AMR-augmented 70 25 6 18 1
0 AMR-only 59 22 2 14 0
3 AMR-only 55 18 2 12 0
Mistral 0 Context-only 76 36 12 27 5
3 Context-only 80 47 22 38 12
0 AMR-augmented 77 38 14 29 6
3 AMR-augmented 77 45 20 36 11
0 AMR-only 66 30 6 21 1
3 AMR-only 66 34 8 25 2
@@
Llama3.1 0 Context-only 47 8
3 Context-only 56 52
0 AMR-augmented 42 15
3 AMR-augmented 56 54
0 AMR-only 37 12
3 AMR-only 44 39
Phi3 0 Context-only 77 27
3 Context-only 82 80
5 Context-only 83 82
0 AMR-augmented 67 39
3 AMR-augmented 79 79
5 AMR-augmented 76 76
0 AMR-only 50 25
3 AMR-only 52 52
5 AMR-only 55 55
Mistral 0 Context-only 48 50
3 Context-only 52 54
0 AMR-augmented 49 50
3 AMR-augmented 34 20
0 AMR-only 33 20
3 AMR-only 33 12